OpenStudy (anonymous):

PLZ PLZ HELP!~ MEDALL!! How many solutions does the system of equations have? y - 7x = -14 and 7y - 49x = -2

OpenStudy (anonymous):

a. one b. two c. infinitely many d. none

OpenStudy (anonymous):

y= 7x-14 and y = 7x -2/7 So 7x-14 = 7x - 2/7 -14 = -2/7 So this has no solutions

OpenStudy (anonymous):

none

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

Thank you!!